Buffalo Company purchased Machine #201 on May 1, 2017. The following information relating to Machine #201 was gathered at the end of May.

Price $ 122,400
Credit terms 2/10, n/30
Freight-in $  1,152
Preparation and installation costs $  5,472
Labor costs during regular production operations $ 15,120


It is expected that the machine could be used for  10 years, after which the salvage value would be zero. Buffalo intends to use the machine for only  8 years, however, after which it expects to be able to sell it for $ 2,160. The invoice for Machine #201 was paid May 5, 2017. Buffalo uses the calendar year as the basis for the preparation of financial statements.

Compute the depreciation expense for the years indicated using the following methods.

(1) Straight-line method for 2017

$

(2) Sum-of-the-years'-digits method for 2018

$

(3) Double-declining-balance method for 2017

$
Answer #1

Calculate cost of machine

Price 122400
Less: Discount (122400*2%) -2448
Freight in 1152
Preparation and installation costs 5472
Cost of machine 126576

1) Straight line method for 2017 = (126576-2160/8)*8/12 = $10368

2) Sum of the years digit method for 2018 :

Sum of year digit = 8+7+6+5+4+3+2+1 = 36

2018 Dep = (126576-2160*8/36)*4/12+(126576-2160*7/36)*8/12 = $25344

3) Double decline rate = 100/8*2 = 25%

2017 Dep = 126576*25%*8/12 = $21096